Using, C = λν E = hν C = 3.00 x 108 m/s h = 6.626 2 x 10-34 J-s (or J/Hz) 8. What is the wavelength of a 1.528 x 10-13 J wave?
Answer #1

Given:

E = 1.528*10^-13 J
use:
E = h*c/lambda
1.528*10^-13 J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda
lambda = 1.301*10^-12 m
Answer: 1.30*10^-12 m